3. La Besurta

RENCLLUSA REFUGE AND IBONS (or PADERNA IBONS) 1 hour

Follow the well-traced track that sets off from the Hospital de Benasque. This will take you to the Rencllusa refuge in 35 minutes. From there take the track that crosses the Maladeta torrent through a wooden bridge. Go uphill on the right of the Alba torrent (the last one on the right of all that go up to the Rencllusa). After 55 minutes you will get to a prairie where the torrent creates some curves. This happens because the first ibón is about to become a prairie. If you jump to the left, you will find the Rencllusa ibón (2,270 m. high), from where you will surely appreciate the magnificent views.